OpenStudy (prince1342):

The first five terms of a sequence are shown below. 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 5n - 1; n ≥ 1 f(n) = 5n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (prince1342):

Giving a medal..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the letter D

OpenStudy (anonymous):

that wht I think

OpenStudy (prince1342):

Another.... A sequence has its first term equal to 8 and each term of the sequence is obtained by adding 6 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 6 and f(n) = f(n - 1) + 8; n > 1 f(1) = 8 and f(n) = f(n - 1) + 6; n > 1 f(1) = 8 and f(n) = f(n - 1) + 6n; n > 1 f(1) = 6 and f(n) = f(n - 1) + 8n; n > 1

OpenStudy (prince1342):

huh

OpenStudy (anonymous):

adding 6 to previous term. previous term f(n-1) calculate f(n)

OpenStudy (prince1342):

Sorry im still confused

OpenStudy (prince1342):

So what would the answer be?

OpenStudy (anonymous):

just add6

OpenStudy (prince1342):

Add 6 to what

OpenStudy (anonymous):

f(1) meansfirst term

OpenStudy (prince1342):

f(7)?

OpenStudy (anonymous):

f(1)=first term =8 f(n)=previous term +6=f(n-1)+6

OpenStudy (prince1342):

So what amswer would that be

OpenStudy (prince1342):

Answer**

OpenStudy (anonymous):

you can see, it is second.

OpenStudy (prince1342):

B?

OpenStudy (anonymous):

yes

OpenStudy (prince1342):

Oooh thxxxxx

OpenStudy (anonymous):

yw